OpenStudy (anonymous):

Please help me. Which of the following forms the correct proportion with 3/15? Answers, 3/15 1/5 3/1

OpenStudy (whpalmer4):

Can you multiply or divide the top and bottom of 3/15 by any single number and get one of the other fractions? An example: 2/4 is the same proportion as 1/2 or 4/8 or 6/12. Multiplying 2/4 by 2 gives you 4/8. Multiplying it by 3 gives you 6/12. Dividing it by 2 gives you 1/2. Therefore, all of those fractions have the same proportion.

OpenStudy (anonymous):

It can only be one answer tho

OpenStudy (whpalmer4):

There are two correct answers in the list for your problem, but let's assume that they want you to find a different fraction which has the same proportion. That leaves you with 1/5 and 3/1 from which to choose. There are infinitely many fractions that have the correct proportion with 3/15. My example was intended to show you that there could be many different answers if you didn't have to select from a list.

OpenStudy (whpalmer4):

So, we'll strike 3/15 from consideration as a possible answer. Is there a number that you can multiply both top and bottom of 1/5 to give 3/15? If so, 1/5 is the correct answer. Otherwise, is there a number you can multiply both top and bottom of 3/1 to give 3/15?

OpenStudy (anonymous):

I cant find a number that multiplys top and bottom of 1/5 to equal 3/15 did you?

OpenStudy (whpalmer4):

What about 3?

OpenStudy (whpalmer4):

\[\frac{ 3 }{ 3 }*\frac{1}{5} = \frac{3*1}{3*5} = \frac{ 3}{15}\]

OpenStudy (anonymous):

oh lol i get it now. thanks for helping